Brian made a burger. The diameter of the burger was 6 cm. What was the area of the burger?


1
Expert's answer
2021-03-30T07:22:31-0400

Since the diameter of the burger was "6\ cm" , the radius of the burger "R" was equal to "3\ cm". It follows that the area of the burger was "\pi R^2=9\pi\approx 28.26\ cm^2."